In comparing the resistance of two coils P and Q with a slide wire bridge,a balance point is obtained when the sliding contact is 30 cm from the zero end of the wire.The resistance P and Q are interchanged and balance is obtained at 120 cm from the same end.Find the ratio of resistance P and Q and the length of the bridge wire.

Grade:12

1 Answers

Princess Insha
6 years ago
As we know slide wire bridge works on h eprinciple of Wheatstone bridge. Where P/Q = L/Total lengthof wire-L ( where L is the length of wire giving balancing position from zero end)Let us asume that total length of wire is x.Then in first case P/Q = 30/x-30 ... (1)In second case Q/P = 120/x-120Implies P/Q = x- 120/120.. (2).From (1) & (2) we have,30/ x - 30 = x - 120 / 120On solving we get a quadratic equation,And value of X comes out to be 150